A real estate agent wants to estimate the average age of those buying investment property in his area. He randomly selects 15 of his clients who purchased an investment property and obtains the data shown. Assume that sequals=10.5. Use this information to answer the following question.

53

62

32

28

55

47

46

37

49

47

32

63

43

38

44

The buyer ages are approximately normally distributed and the sample does not contain any outliers. Construct a 90​% confidence interval for the mean age for all the real estate​ agent's clients who purchased investment property.
Answer #1

\sample mean, xbar = 45.07
sample standard deviation, s = 10.5
sample size, n = 15
degrees of freedom, df = n - 1 = 14

Given CI level is 90%, hence α = 1 - 0.9 = 0.1
α/2 = 0.1/2 = 0.05, tc = t(α/2, df) = 1.761

CI = (xbar - tc * s/sqrt(n) , xbar + tc * s/sqrt(n))
CI = (45.07 - 1.761 * 10.5/sqrt(15) , 45.07 + 1.761 * 10.5/sqrt(15))
CI = (40.30 , 49.84)